OpenStudy (anonymous):

Find a polynomial function of degree 3 with -2, 2, 5 as zeros.

hero (hero):

If -2, 2, and 5 are zeroes then x = 2 x = -2 x = 5 Move everything to the left. Then x - 2 = 0 x + 2 = 0 x - 5 = 0 By zero product property (x - 2)(x + 2)(x - 5) = 0 Expand the left side to form a cubic polynomial, then set it equal to f(x).